# Sylva: Tailoring Personalized Adversarial Defense in Pre-trained Models via Collaborative Fine-tuning

### v1

@@ 4.5. Ablation Experiments (RQ4)
$\gamma$ 0.9 55.03 59.03 $\beta$ 20 56.87 57.27
0.7 54.65 58.35 10 56.22 58.34
0.5 54.70 57.25 5 55.03 59.03
0.3 52.97 56.24 1 51.29 61.25
$\epsilon$ 0.9 55.03 59.03 $B$ 50 49.24 53.98
0.7 55.01 58.93 100 53.04 56.37
0.5 54.31 58.35 300 55.03 59.03
0.3 55.25 58.22 500 55.39 60.77
$r$ 8 54.86 58.63 PGD Strength 3 48.04 61.26
4 55.03 59.03 5 52.16 60.64
2 55.42 58.36 10 55.03 59.03
1 53.26 57.28 15 56.96 56.89
4.5.2. Impact of Different Hyperparameters.
To investigate the impact of hyperparameters on Sylva, we conduct a comprehensive set of experiments on the ViT-B model using the CIFAR-10 dataset. This analysis evaluates the effects of key hyperparameters, including $\gamma$ and $\epsilon$ in the Phase 1 loss function, $\beta$ and $B$ in Phase 2, the hidden space dimension $r$ for LoRA, and the PGD strength employed in adversarial training. Table 5 presents the corresponding results for adversarial robustness and benign accuracy, providing a detailed comparison across different hyperparameter settings. The findings demonstrate that the default hyperparameter settings deliver strong overall performance, balancing robustness and accuracy effectively. Notably, smaller sampling sizes ($B$) in Phase 2 result in significantly worse outcomes, highlighting the importance of an adequate sampling strategy. In contrast, larger values of $B$ show diminishing returns, with minimal performance gains and increased training time. Furthermore, the hidden space dimension $r$ for LoRA shows minimal impact on the final results; even with $r=1$, performance remains excellent. This observation aligns with LoRA’s established efficiency and adaptability in traditional fine-tuning settings, further validating its applicability in adversarial training scenarios. Additionally, we evaluate the effects of these hyperparameters on three other datasets, as well as on scenarios with varying numbers of clients. The detailed results are provided in Appendix E.
Answer to RQ4: Sylva excels in balancing robustness and accuracy through its modular design, efficient aggregation, and adaptive hyperparameter tuning, validated across diverse datasets and scenarios.
5. Discussion
This section explores four key directions for further improving Sylva, focusing on supporting heterogeneous deployment across devices with varying computational capacities, enhancing experimental design with larger-scale models, refining the threat model to include training-time attacks, and broadening its applicability to multi-modal training and diverse adversarial scenarios.
Adding Heterogeneous Model Support for Device Adaptation. The rapid advancement of AI models has driven industries to adopt increasingly powerful models for their applications. However, in industries such as the automotive sector, the long operational lifespan of vehicles (typically around 18 years (ASME, 2025)) leads to substantial variability in hardware capabilities across the system. As a result, not all vehicles are able to support the latest, most demanding models. To maximize performance, each vehicle should deploy the best model that its hardware can reliably support. Devices with varying computational capacities may require different model sizes, posing a challenge for maintaining the effectiveness of heterogeneous LoRA models fine-tuned via Sylva. Future work could explore methods for handling heterogeneous models within Sylva’s adversarial training, enhancing adaptability across diverse devices through dynamic scaling or pruning.
Expanding experiments with larger-scale models. Our algorithm has shown strong performance on traditional open-source ViT models. However, as model sizes continue to grow with the development of large-scale architectures, the robustness of massive vision models, such as LLaVA (Liu et al., 2023) and Qwen-VL (Bai et al., 2023) with hundreds of billions of parameters, remains an open question. Leveraging LoRA for fine-tuning, Sylva can be seamlessly integrated into these models, extending its applicability to unprecedented scales. While such experiments are resource-intensive, they represent a feasible and valuable direction for future exploration.
Strengthening the threat model. The current threat model assumes attacks occur during the inference phase in federated adversarial training. However, vulnerabilities may arise during training, even with secure devices, exposing models to potential risks. Addressing the detection and elimination of malicious clients during distributed training is a critical challenge. Integrating Byzantine defense techniques into Sylva’s Phase 1 offers a promising solution (Ozdayi et al., 2021; Shejwalkar and Houmansadr, 2021; Fang et al., 2020; Ma et al., 2022), paving the way for future research to further enhance model robustness in this direction.
Adapting to multi-modal training and attacks. With the rapid evolution of large models in text (Hoffmann et al., 2022) and multi-modal (Radford et al., 2021) applications, adversarial attack and defense algorithms are increasingly being developed to address their unique vulnerabilities (Liu et al., 2020; Zhou et al., 2023). Integrating these defense mechanisms into Sylva can further enhance its adaptability, enabling efficient and task-specific adversarial training across diverse downstream applications. This integration represents a promising direction for expanding Sylva’s applicability and effectiveness.
6. Related works
6.1. Adversarial Attack
Recent studies have focused on generating adversarial examples through subtle perturbations leading to misclassifications. Goodfellow et al. proposed FGSM (Goodfellow et al., 2014), highlighting neural networks’ vulnerability due to their linearity and introducing a simple adversarial training approach. Tramèr et al. improved robustness against black-box attacks with RFGSM (Tramèr et al., 2018), which integrates perturbations from multiple models. Madry et al. introduced PGD (Madry et al., 2018), providing robust optimization with security guarantees. Liu et al. developed EOTPGD (Liu et al., 2018), a Bayesian framework incorporating randomness for greater robustness. Croce et al. proposed APGD (Croce and Hein, 2020), a parameter-free ensemble attack for evaluating defense strategies. Su and Modas et al. introduced the One Pixel Attack (Su et al., 2019) and Sparsefool (Modas et al., 2019), optimizing minimal pixel changes to reduce computational overhead. Ban et al. presented PAP (Ban and Dong, 2022), leveraging Low-Level Layer Lifting Attacks (L4A) to manipulate neuron activations in pre-trained models’ lower layers. Zhou et al. introduced AdvCLIP (Zhou et al., 2023), generating downstream-agnostic adversarial examples with cross-modal pre-trained encoders.
6.2. Adversarial Defence in Federated Learning
Federated learning (FL) enables distributed model training while preserving user privacy (McMahan et al., 2017; Zhan et al., 2020; Liu et al., 2022c), but faces challenges due to statistical heterogeneity in device data distributions (Zhang et al., 2021; Qi et al., 2023; Luo et al., 2021). To overcome the limitations of global models in non-IID scenarios, personalized federated learning approaches have been proposed to train client-specific models (Wang et al., 2024; Zhang et al., 2023a). Many works leverage parameter-efficient methods in FL for LLMs, including adapter tuning (Qi et al., 2024), prompt learning (Guo et al., 2023), and localized adjustments (Sun et al., 2024). With the continuous development of edge computing, many works have implemented adversarial attacks and defenses based on the FL framework (Krauß and Dmitrienko, 2023; Pasquini et al., 2022; Yazdinejad et al., 2024). Zizzo et al. investigated federated adversarial training, comparing model performance in both centralized and distributed FL settings (Zizzo et al., 2020). Li et al. provided a convergence proof for this federated adversarial learning approach (Li et al., 2023b). Zhou et al. addressed the aggregation error in FL, breaking it down into bias and variance (Zhou et al., 2022). Additionally, Chen et al. explored certified defenses against adversarial examples in FL, providing an alternative direction for enhancing robustness (Chen et al., 2022). Zhang et al. proposed DBFAT (Zhang et al., 2023b), which adjusts sample weights based on PGD steps for generating adversarial examples. Hong et al. introduced a strategy propagating adversarial robustness from resource-rich to resource-poor users via batch normalization (Hong et al., 2023). Qi et al. developed FairVFL, a privacy-preserving framework enhancing fairness in vertical federated learning through adversarial learning (Qi et al., 2022).
7. Conclusion
In this paper, we introduce Sylva, the first adversarial defense algorithm tailored for pre-trained models in distributed settings. Sylva delivers robust and accurate models even with imbalanced data distributions. Through preliminary experiments, we identify key challenges and highlight the advantages of efficient parameter fine-tuning. Building on this, we propose a two-phase training framework: Phase 1 utilizes a novel loss function and personalized federated fine-tuning for enhanced robustness, while Phase 2 applies a game-based layer freezing strategy to balance robustness and accuracy. Experiments demonstrate that Sylva not only outperforms state-of-the-art adversarial defenses in terms of robustness and accuracy but also does so with minimal computational overhead, making it suitable for edge deployment.
Appendix A Details of Experimental Setup
A.1. Datasets and Models
  • CIFAR-10 (Krizhevsky et al., 2009): A dataset of 60,000 32$\times$32 color images across 10 classes, with 50,000 for training and 10,000 for testing.
  • STL-10 (Coates et al., 2011): A dataset with 10 object classes, consisting of 5,000 training and 8,000 test images at 96$\times$96 resolution, commonly used for unsupervised and semi-supervised learning benchmarks.
  • GTSRB (Stallkamp et al., 2011): A traffic sign recognition dataset with over 50,000 images spanning 43 classes, serving as a benchmark for intelligent transportation systems.
  • CIFAR-100 (Krizhevsky et al., 2009): Similar to CIFAR-10 but with 100 classes, comprising 60,000 images (600 per class), offering a greater challenge in image classification.
The Vision Transformer (ViT) is a transformer-based model for image classification, treating images as sequences of patches. We use three pre-trained variants: ViT-Tiny (ViT-T/16), ViT-Base (ViT-B/16), and ViT-Large (ViT-L/16), differing in scale and complexity. Their configurations, summarized in Table 6, enable performance evaluation under varying computational demands.
A.2. Attacks
Building on the threat model outlined in Section 2, attacker methods can be broadly divided into white-box and gray-box attacks. In our experiments, we focus on four representative attack methods, detailed as follows:
  • FGSM (Goodfellow et al., 2014): A white-box attack that perturbs inputs along the gradient direction of the loss function, efficiently generating small but impactful adversarial examples.
  • PGD (Madry et al., 2018): An iterative extension of FGSM, applying multiple small perturbations projected back into the allowed space for a more thorough adversarial optimization.
  • SparseFool (Modas et al., 2019): Focuses on creating sparse perturbations by minimizing pixel changes, offering effective attacks with high computational efficiency for high-dimensional data.
  • PAP (Ban and Dong, 2022): A gray-box attack targeting pre-trained models, effective even against fine-tuned models by leveraging low-level layer manipulations and noise augmentation for strong transferability.
Table 6. Details of vision transformer models
Model Layers Hidden Size Intermediate Size Heads
ViT-T/16 12 192 768 3
ViT-B/16 12 768 3072 12
ViT-L/16 24 1024 4096 16
Figure 12. Visualization under different attack algorithms
A.3. Baselines
To demonstrate the superiority of Sylva, we first examine commonly used and novel defense algorithms. For local defense, the following three algorithms are employed:
  • PGT-AT (Madry et al., 2018): PGD-AT employs the PGD algorithm to generate adversarial examples during training, thereby enhancing model robustness. By repeatedly exposing the model to adversarial perturbations, it improves the model’s ability to resist various attacks effectively.
  • TRADES (Zhang et al., 2019): TRADES achieves a balance between accuracy on benign data and robustness against adversarial attacks. It minimizes a carefully designed loss function that combines the standard classification loss with a penalty term measuring output discrepancies between benign and adversarial examples.
  • Gen-AF (Zhou et al., 2024): Gen-AF improves model robustness through a two-stage fine-tuning approach for pre-trained models. It first optimizes the encoder using genetic regularization, followed by the selection and refinement of robust layers, leading to enhanced accuracy and robustness across diverse datasets.
  • FedAvg (McMahan et al., 2017): FedAvg is a federated learning algorithm where each client trains a local model and shares updates with the server, which computes a weighted average to update the global model.
  • FedProx (Li et al., 2020b): FedProx extends FedAvg by adding a proximal term to the loss function, mitigating the impact of heterogeneous data across clients. This regularization ensures that local models remain close to the global model, improving convergence and robustness.
  • DBFAT (Zhang et al., 2023b): DBFAT enhances adversarial robustness in federated learning through local re-weighting and global regularization, improving both accuracy and robustness, especially in non-IID settings.
  • Per-Adv (Allouah et al., 2024): Per-Adv enhances robustness against Byzantine attacks by introducing a personalized loss function, wherein each client interacts with others to collaboratively construct a robust interpolated objective.
  • Per-LoRA (Wagner et al., 2024): Per-LoRA leverages LoRA-based fine-tuning to perform local updates of large models, while additionally computing trust weights from each client’s LoRA parameters, which are used to guide the personalized update of the local model.
Specifically, we integrate the FedAvg and FedProx algorithms with the three aforementioned local defense algorithms, yielding the following configurations: FA-PGD-AT, FA-TRADES, and FA-Gen-AF, as well as FP-PGD-AT, FP-TRADES, and FP-Gen-AF. These combinations leverage the strengths of both federated optimization strategies and local defense mechanisms to enhance model robustness. For the two personalized algorithms, Per-Adv and Per-LoRA, we apply appropriate optimizations to adapt them to our specific tasks. We utilize the TRADES loss function for both algorithms to facilitate adversarial training. In the case of Per-Adv, we assume an adversary without Byzantine attacks to better align with the threat model outlined in this study. For Per-LoRA, we employ strategy 2, as proposed by the authors, to compute the trust weights.
(a) IID
(b) Dir-10
(c) Dir-6
(d) Dir-3
Figure 13. Visualization of Dirichlet distribution
Table 7. Performance comparison of Sylva and baseline under different datasets and attack algorithms (ViT-T/16)
Datasets Baselines Benign(BA) FGSM PGD SparseFool PAP
CIFAR-10 FA-PGD-AT 37.54 41.82 40.95 46.10 42.23
FA-TRADES 48.24 43.10 42.58 44.15 41.48
FA-Gen-AF 50.12 44.68 42.74 43.59 42.61
FP-PGD-AT 39.91 42.35 41.62 45.86 42.48
FP-TRADES 49.58 44.90 42.69 43.53 42.10
FP-Gen-AF 51.73 45.50 43.91 45.57 44.91
DBFAT 51.45 43.62 42.80 45.35 44.50
Per-Adv 39.96 41.93 41.32 46.50 42.77
Per-LoRA 50.89 45.90 43.03 45.21 44.30
Sylva(Ours) 55.92 50.02 52.45 50.25 49.58
STL-10 FA-PGD-AT 32.81 36.12 34.89 38.45 37.67
FA-TRADES 40.31 35.56 34.88 37.91 37.40
FA-Gen-AF 42.29 36.85 35.30 37.91 36.77
FP-PGD-AT 34.41 36.48 35.21 39.35 37.89
FP-TRADES 41.10 37.21 35.22 37.00 36.88
FP-Gen-AF 43.10 38.92 36.67 39.60 38.53
DBFAT 42.75 36.77 35.48 39.14 38.11
Per-Adv 32.92 36.83 35.01 38.25 38.01
Per-LoRA 43.06 38.90 36.82 39.25 38.44
Sylva(Ours) 48.21 42.37 41.12 43.82 42.78
GTSRB FA-PGD-AT 55.73 63.52 64.25 69.12 68.22
FA-TRADES 71.15 64.10 65.28 67.39 68.52
FA-Gen-AF 74.85 65.49 66.10 67.50 68.10
FP-PGD-AT 59.15 64.38 65.45 68.23 67.90
FP-TRADES 72.30 65.75 66.52 67.37 68.20
FP-Gen-AF 75.92 66.10 66.48 67.75 69.71
DBFAT 75.40 65.38 65.53 66.75 68.95
Per-Adv 60.73 63.06 64.25 67.21 67.53
Per-LoRA 74.33 65.80 65.71 67.05 68.32
Sylva(Ours) 76.90 68.75 68.90 69.93 69.03
CIFAR-100 FA-PGD-AT 11.89 24.32 23.75 25.92 24.70
FA-TRADES 22.60 23.52 22.11 23.77 23.91
FA-Gen-AF 22.39 24.00 23.12 24.30 23.68
FP-PGD-AT 14.02 24.45 24.18 26.12 24.90
FP-TRADES 23.41 22.78 23.75 25.18 24.05
FP-Gen-AF 23.18 24.10 23.99 25.95 25.30
DBFAT 23.28 23.89 23.98 25.50 26.88
Per-Adv 14.65 23.87 22.64 23.85 24.04
Per-LoRA 23.27 24.20 23.75 24.89 25.06
Sylva(Ours) 24.92 25.12 24.82 26.25 27.15
Table 8. Performance comparison of Sylva and baseline under different datasets and attack algorithms (ViT-L/16)
Datasets Baselines Benign(BA) FGSM PGD SparseFool PAP
CIFAR-10 FA-PGD-AT 40.50 47.21 43.89 49.45 45.58
FA-TRADES 52.37 46.18 44.47 47.00 44.64
FA-Gen-AF 54.75 48.17 46.87 47.27 46.78
FP-PGD-AT 42.14 48.62 45.09 50.02 46.87
FP-TRADES 54.17 48.02 46.68 48.78 46.56
FP-Gen-AF 56.07 49.94 47.83 50.10 48.48
DBFAT 54.92 47.01 46.13 50.20 49.72
Per-Adv 41.34 47.35 44.02 48.35 45.66
Per-LoRA 55.83 48.92 47.60 49.57 48.21
Sylva(Ours) 61.51 54.01 56.39 54.12 53.26
STL-10 FA-PGD-AT 35.89 41.05 39.88 41.95 41.05
FA-TRADES 43.70 39.15 38.11 41.80 40.93
FA-Gen-AF 46.00 40.48 39.12 41.98 41.85
FP-PGD-AT 37.78 41.77 39.14 42.51 41.69
FP-TRADES 45.02 40.72 39.12 40.93 40.99
FP-Gen-AF 47.20 41.03 40.49 42.12 43.02
DBFAT 46.79 42.17 39.91 41.14 42.78
Per-Adv 37.95 41.23 39.48 41.52 41.12
Per-LoRA 46.95 41.78 41.20 42.05 42.11
Sylva(Ours) 50.51 45.01 43.52 46.47 45.42
GTSRB FA-PGD-AT 61.69 69.99 70.69 73.65 73.48
FA-TRADES 75.24 69.58 70.40 72.34 74.15
FA-Gen-AF 77.86 70.63 69.92 74.12 73.36
FP-PGD-AT 63.98 71.57 70.66 73.62 72.71
FP-TRADES 78.04 71.58 71.62 72.65 72.76
FP-Gen-AF 81.14 72.81 71.83 73.95 75.07
DBFAT 80.96 70.77 70.88 72.47 74.30
Per-Adv 65.52 69.29 69.37 72.55 73.18
Per-LoRA 78.83 71.52 70.25 72.36 74.01
Sylva(Ours) 82.51 74.39 74.20 75.64 75.84
CIFAR-100 FA-PGD-AT 14.35 26.59 26.04 28.29 27.33
FA-TRADES 24.73 25.50 24.93 25.95 26.14
FA-Gen-AF 24.23 26.05 25.81 26.81 26.09
FP-PGD-AT 15.82 26.47 26.04 28.86 27.46
FP-TRADES 25.32 24.42 26.05 27.32 26.33
FP-Gen-AF 25.70 26.11 27.89 28.15 27.30
DBFAT 25.66 25.78 26.91 27.79 29.04
Per-Adv 16.80 25.88 26.01 26.72 27.04
Per-LoRA 25.97 25.99 26.32 27.85 28.03
Sylva(Ours) 29.62 28.33 28.87 29.77 29.44
@@ A.3. Baselines
Figure 14. Comparison of adversarial training performance under different heterogeneity levels (ViT-T/16)
(a) CIFAR-10
(b) STL-10
(c) GTSRB
(d) CIFAR-100
Figure 15. Comparison of adversarial training performance under different heterogeneity levels (ViT-L/16)
A.4. Data Heterogeneity
We adopt the widely used data heterogeneity partitioning method based on the Dirichlet distribution (Li et al., 2020a), which effectively captures the non-IID nature of data distributions typically observed in real-world scenarios. For the CIFAR-10 dataset, we distribute the data among 15 clients with varying Dirichlet distribution parameters, ensuring that both the training and test sets for each client follow the same distribution. This method realistically simulates data heterogeneity, providing a robust setup for evaluating federated learning algorithms in non-IID environments.
(a) ViT-T
(b) ViT-B
(c) ViT-L
Figure 16. Relationship between the number of clients and communication time
Table 2 provides a comprehensive comparison of Sylva against baseline models utilizing the ViT-B architecture across multiple datasets. The results emphasize Sylva’s superior performance in terms of both adversarial robustness and benign accuracy under a wide range of adversarial attack scenarios. Notably, Sylva consistently surpasses the baseline methods, demonstrating its capability to effectively manage diverse attack conditions while maintaining exceptional levels of adversarial reliability.
To further evaluate the adaptability of Sylva, we extend our analysis to models of varying sizes, specifically ViT-T and ViT-L, with detailed results presented in Table 7 and Table 8. The findings highlight that Sylva maintains strong defense capabilities across different model scales, delivering robust performance even as model complexity changes. Sylva outperforms both traditional distributed algorithms and personalized algorithms, consistently demonstrating superior performance in terms of both robustness and adaptability. Notably, while the defense capabilities of baseline methods often degrade significantly with smaller models, such as ViT-T, Sylva demonstrates exceptional adaptability. It consistently achieves an effective balance between adversarial robustness and benign accuracy, underscoring its versatility and resilience under varying conditions. These findings highlight the remarkable versatility of Sylva, demonstrating its suitability for a broad spectrum of model architectures and adversarial scenarios. Its consistent ability to deliver robust performance across diverse datasets and varying model scales underscores its potential as a dependable solution for adversarial training. This adaptability makes Sylva an effective choice for applications ranging from simple to highly complex environments, further cementing its value in enhancing model resilience.
(a) GTSRB
(b) CIFAR-100
Figure 17. Impact of different loss modules on adversarial robustness and benign accuracy on GTSRB and CIFAR-100
In Section 4.3, we thoroughly analyze the adversarial defense performance of the ViT-B model under diverse heterogeneity conditions, with Sylva demonstrating remarkable defense capabilities and robustness throughout. To gain deeper insights into how models of varying scales respond to heterogeneity, we extend the evaluation to the ViT-T and ViT-L models, systematically assessing their defense performance under different levels of heterogeneity, as depicted in Fig. 14 and Fig. 15, respectively.
The results demonstrate that Sylva consistently outperforms its counterparts across various model sizes, effectively maintaining both accuracy and robustness, even in highly challenging heterogeneous environments where data distribution and computational resources vary significantly. Notably, smaller models, such as ViT-T, exhibit more rapid convergence during training, enabling them to achieve relatively stable performance despite the increasing heterogeneity of the data. This finding underscores Sylva’s adaptability in managing the intricate balance between model complexity and environmental variability. Moreover, compared to the other two personalized baseline algorithms, the smaller performance fluctuations observed under heightened heterogeneity indicate that Sylva’s personalized framework effectively mitigates the detrimental impacts of data heterogeneity by tailoring defense strategies to individual clients. These findings highlight the exceptional versatility of Sylva, making it an ideal choice for adversarial training across systems characterized by diverse heterogeneity and a broad spectrum of model sizes.
(a) GTSRB
(b) CIFAR-100
Figure 18. Impact of ball-tree-based aggregation algorithm on GTSRB and CIFAR-100
(a) GTSRB
(b) CIFAR-100
Figure 19. Impact of different phases on the trade-off between robustness and accuracy for GTSRB and CIFAR-100
Table 9. The Impact of Hyperparameters in STL-10
Para Value AR BA Para Value AR BA
$\gamma$ 0.9 41.48 49.11 $\beta$ 20 42.97 46.81
0.7 41.26 48.83 10 42.50 47.25
0.5 40.55 48.80 5 41.48 49.11
0.3 40.12 48.05 1 40.79 50.02
$\epsilon$ 0.9 41.48 49.11 $B$ 50 39.27 48.05
0.7 40.96 49.05 100 41.10 47.92
0.5 40.23 49.24 300 41.48 49.11
0.3 40.57 48.76 500 41.76 50.21
$r$ 8 41.72 49.34 PGD Strength 3 40.02 52.17
4 41.48 49.11 5 41.07 50.81
2 41.25 49.03 10 41.48 49.11
1 40.27 48.65 20 42.52 47.33
In Section 4.4, we emphasize the efficiency advantages of Sylva in real-world systems composed of diverse edge devices operating under adversarial training scenarios. To delve deeper into its communication efficiency, we perform additional experiments to assess the communication overhead in distributed systems of varying scales. By leveraging the edge devices described in Section 4.4, we simulate distributed systems with 5, 10, 15, and 20 clients, each performing adversarial training while concurrently communicating with the server. As the client count rises, network congestion emerges as a critical challenge, potentially undermining the timeliness and reliability of adversarial updates.
Fig. 16 presents a comparative analysis of communication time for distributed training across three distinct models under different client counts. The results clearly demonstrate a proportional increase in communication time as the number of clients grows, primarily due to heightened network congestion. Crucially, Sylva and Per-LoRA exhibit a significant advantage over traditional algorithms, particularly with larger models, by transmitting only the lightweight LoRA parameters. Since Sylva only transmits the backbone’s LoRA parameters and keeps the classifier local to preserve personalization, it incurs lower communication overhead, making it more efficient for real-world deployments. This approach drastically reduces communication overhead, underscoring Sylva’s practicality and efficiency for deployment in real-world distributed systems. Its ability to mitigate communication challenges is especially valuable in scenarios involving large-scale models and high client densities, further cementing its suitability for such environments.
Table 10. The Impact of Hyperparameters in GTSRB
Para Value AR BA Para Value AR BA
$\gamma$ 0.9 72.40 80.49 $\beta$ 20 72.14 78.77
0.7 72.34 79.25 10 72.80 80.04
0.5 72.04 78.95 5 72.40 80.49
0.3 71.83 78.24 1 70.92 81.94
$\epsilon$ 0.9 72.40 80.49 $B$ 50 70.20 78.04
0.7 72.37 78.37 100 71.29 80.82
0.5 71.02 79.42 300 72.40 80.49
0.3 71.25 78.10 500 72.43 79.95
$r$ 8 72.66 79.42 PGD Strength 3 70.94 82.03
4 72.40 80.49 5 71.02 81.72
2 72.98 80.93 10 72.40 80.49
1 71.62 78.39 20 73.97 77.92
Table 11. The Impact of Hyperparameters in CIFAR-100
Para Value AR BA Para Value AR BA
$\gamma$ 0.9 26.23 26.95 $\beta$ 20 26.54 25.10
0.7 26.05 26.55 10 26.40 25.15
0.5 26.09 26.72 5 26.23 26.95
0.3 25.98 25.76 1 24.88 28.24
$\epsilon$ 0.9 26.23 26.95 $B$ 50 25.87 25.90
0.7 26.42 25.87 100 26.05 26.45
0.5 25.87 25.58 300 26.23 26.95
0.3 25.90 25.07 500 26.20 26.97
$r$ 8 26.44 27.01 PGD Strength 3 25.04 28.03
4 26.23 26.95 5 25.88 27.89
2 26.30 27.04 10 26.23 26.95
1 26.03 26.44 20 27.12 24.87
Appendix E Additional Ablation Study Resultsn (RQ4)
In Section 4.5, we evaluate the impact of Sylva’s key modules on adversarial training models, focusing on CIFAR-10 and STL-10 datasets. This includes analyzing the loss function design, aggregation methods, and two-phase training framework, all critical to Sylva’s performance. To validate these findings, we extend experiments to GTSRB and CIFAR-100 datasets. The results, presented in Fig. 17, Fig. 18, and Fig. 19, confirm the consistent contributions of each module. These trends align with the main text, demonstrating Sylva’s robustness and strong generalization across datasets of varying complexity, highlighting its versatility in adversarial training scenarios.
We evaluate the impact of various hyperparameters, with detailed results shown in Table 9, Table 10, and Table 11. These findings align with those on CIFAR-10, revealing consistent trends across all datasets. Specifically, decreases in $\gamma$ and $\epsilon$ lead to declines in both adversarial robustness and benign accuracy, underscoring their importance in maintaining model performance. In Phase 2, reducing $\beta$ slightly improves benign accuracy but significantly reduces adversarial robustness, highlighting the trade-off between these metrics. These results stress the need for careful hyperparameter tuning to balance robustness and accuracy in adversarial training frameworks.
(a) CIFAR-10
(b) STL-10
(c) GTSRB
(d) CIFAR-100
Figure 20. Comparison of adversarial training performance under different number of clients (ViT-B/16)
To comprehensively evaluate the performance of the proposed framework, we conduct experiments under varying client numbers, as illustrated in Fig. 20. For comparison, we select several high-performing baseline algorithms, including FA-Gen-AF, FP-Gen-AF, and DBFAT, to ensure a robust analysis. In the multi-client scenario, the number of clients is systematically varied among 15, 12, 9, and 6, with experiments conducted on four distinct datasets to assess both adversarial robustness and benign accuracy comprehensively. The results reveal a clear trend: as the number of clients decreases, the overall performance improves, likely due to the reduced data heterogeneity associated with fewer clients. Notably, across all configurations, Sylva consistently outperforms the baseline algorithms, demonstrating superior results in both adversarial robustness and benign accuracy. These findings underscore the framework’s effectiveness and adaptability, highlighting its capacity to excel across a diverse range of multi-client settings, irrespective of the dataset or client configuration.

### v3

@@ Appendix C Additional Results on the Performance Under Different Heterogeneity Levels (RQ2)
$\gamma$ 0.9 44.07 52.13 $\beta$ 20 46.25 50.13
0.7 44.30 51.67 10 46.91 50.37
0.5 43.87 51.22 5 44.07 52.13
0.3 43.15 50.48 1 44.75 51.66
$\epsilon$ 0.9 44.07 52.13 $B$ 5% 42.76 51.50
0.7 43.34 51.30 10% 44.07 52.13
0.5 42.52 51.42 20% 44.79 52.26
0.3 42.81 51.77 30% 44.36 53.15
$r$ 8 44.89 52.53 PGD Strength 3 43.14 55.33
4 44.07 52.13 5 44.53 53.80
2 44.58 53.21 10 44.07 52.13
1 43.77 52.86 20 45.60 50.51
In section IV-D, we emphasize the efficiency advantages of Lorica in real-world systems composed of diverse edge devices operating under adversarial training scenarios. To delve deeper into its communication efficiency, we perform additional experiments to assess the communication overhead in distributed systems of varying scales. By leveraging the edge devices described in section IV-D, we simulate distributed systems with 5, 10, 15, and 20 clients, each performing adversarial training while concurrently communicating with the server. As the client count rises, network congestion emerges as a critical challenge, potentially undermining the timeliness and reliability of adversarial updates.
fig. 18 presents a comparative analysis of communication time for distributed training across three distinct models under different client counts. The results clearly demonstrate a proportional increase in communication time as the number of clients grows, primarily due to heightened network congestion. Crucially, Lorica and Per-LoRA exhibit a significant advantage over traditional algorithms, particularly with larger models, by transmitting only the lightweight LoRA parameters. Since Lorica only transmits the backbone’s LoRA-FA parameters and keeps the classifier local to preserve personalization, it incurs lower communication overhead, making it more efficient for real-world deployments. This approach drastically reduces communication overhead, underscoring Lorica’s practicality and efficiency for deployment in real-world distributed systems. Its ability to mitigate communication challenges is especially valuable in scenarios involving large-scale models and high client densities, further cementing its suitability for such environments.
TABLE X: The Impact of Hyperparameters in GTSRB
Para Value AR BA Para Value AR BA
$\gamma$ 0.9 73.92 81.54 $\beta$ 20 73.21 79.77
0.7 73.33 80.57 10 73.58 81.62
0.5 73.46 80.10 5 73.92 81.54
0.3 72.92 79.59 1 72.39 81.42
$\epsilon$ 0.9 73.92 81.54 $B$ 5% 71.80 80.04
0.7 73.64 79.59 10% 73.92 81.54
0.5 72.30 80.76 20% 73.36 81.38
0.3 72.47 79.35 30% 72.69 80.52
$r$ 8 73.88 80.74 PGD Strength 3 72.08 83.28
4 73.92 81.54 5 72.31 82.52
2 73.54 81.26 10 73.92 81.54
1 72.68 79.67 20 74.71 79.27
TABLE XI: The Impact of Hyperparameters in CIFAR-100
Para Value AR BA Para Value AR BA
$\gamma$ 0.9 27.77 27.93 $\beta$ 20 27.54 26.13
0.7 27.06 27.35 10 27.48 26.10
0.5 27.13 27.44 5 27.77 27.93
0.3 26.15 26.57 1 25.44 29.47
$\epsilon$ 0.9 27.77 27.93 $B$ 5% 26.67 26.58
0.7 27.30 27.02 10% 27.77 27.93
0.5 26.85 26.29 20% 27.33 28.01
0.3 26.27 26.08 30% 27.32 28.12
$r$ 8 27.55 28.13 PGD Strength 3 26.15 29.09
4 27.77 27.93 5 26.71 28.68
2 27.04 28.15 10 27.77 27.93
1 27.01 27.49 20 28.19 26.02
Appendix E Additional Results on the Performance of Byzantine Defence (RQ4)
In section IV-E, we evaluate the defensive performance of Lorica on CIFAR-10 and STL-10. In this subsection, we further extend our analysis to more challenging and complex datasets, namely GTSRB and CIFAR-100. The corresponding results are presented in fig. 19 and fig. 20. The experimental findings demonstrate that, compared with Sylva and its enhanced variants, Lorica consistently achieves superior robustness and accuracy across all evaluated datasets and under diverse attack settings. Notably, even on more fine-grained such as CIFAR-100 and domain-specific such as GTSRB classification tasks, where class diversity and distributional complexity are higher, Lorica maintains stable defensive performance. These results further indicate that Lorica exhibits strong resilience against Byzantine behaviors, effectively mitigating the impact of malicious updates across diverse and complex data distributions.
Appendix F Additional Ablation Study Results (RQ5)
In section IV-F, we evaluate the impact of Lorica’s key modules on adversarial training models, focusing on CIFAR-10 and STL-10 datasets. This includes analyzing the loss function design, aggregation methods, and two-phase training framework, all critical to Lorica’s performance. To validate these findings, we extend experiments to GTSRB and CIFAR-100 datasets. The results, presented in fig. 21, fig. 22, and fig. 23, confirm the consistent contributions of each module. These trends align with the main text, demonstrating Lorica’s robustness and strong generalization across datasets of varying complexity, highlighting its versatility in adversarial training scenarios.
We evaluate the impact of various hyperparameters, with detailed results shown in table IX, table X, and table XI. These findings align with those on CIFAR-10, revealing consistent trends across all datasets. Specifically, decreases in $\gamma$ and $\epsilon$ lead to declines in both adversarial robustness and benign accuracy, underscoring their importance in maintaining model performance. In Phase 2, reducing $\beta$ slightly improves benign accuracy but significantly reduces adversarial robustness, highlighting the trade-off between these metrics. Different training budgets $B$ also exhibit a certain degree of sensitivity in the results. These results stress the need for careful hyperparameter tuning to balance robustness and accuracy in adversarial training frameworks.
@@ Appendix F Additional Ablation Study Results (RQ5)
Figure 24: Comparison of adversarial training performance under different number of clients (ViT-B/16)
To comprehensively evaluate the performance of the proposed framework, we conduct experiments under varying client numbers, as illustrated in fig. 24. For comparison, we select several high-performing baseline algorithms, including FA-Gen-AF, FP-Gen-AF, DBFAT and Sylva, to ensure a robust analysis. In the multi-client scenario, the number of clients is systematically varied among 15, 12, 9, and 6, with experiments conducted on four distinct datasets to assess both adversarial robustness and benign accuracy comprehensively. The results reveal a clear trend: as the number of clients decreases, the overall performance improves, likely due to the reduced data heterogeneity associated with fewer clients. Notably, across all configurations, Lorica consistently outperforms the baseline algorithms, demonstrating superior results in both adversarial robustness and benign accuracy. These findings underscore the framework’s effectiveness and adaptability, highlighting its capacity to excel across a diverse range of multi-client settings, irrespective of the dataset or client configuration.